# Polar Sparsity: High Throughput Batched LLM Inferencing with Scalable Contextual Sparsity

### v2

@@ Abstract
Accelerating large language model (LLM) inference is critical for real-world deployments requiring high throughput and low latency. Contextual sparsity, where each token dynamically activates only a small subset of the model parameters, shows promise but does not scale to large batch sizes due to union of active neurons quickly approaching dense computation. We introduce Polar Sparsity, highlighting a key shift in sparsity importance from MLP to Attention layers as we scale batch size and sequence length. While MLP layers become more compute-efficient under batching, their sparsity vanishes. In contrast, attention becomes increasingly more expensive at scale, while their head sparsity remains stable and batch-invariant. We develop hardware-efficient, sparsity-aware GPU kernels for selective MLP and Attention computations, delivering up to $\times2$ end-to-end speedups for models like OPT, LLaMA-2 & 3, across various batch sizes and sequence lengths without compromising accuracy. To our knowledge, this is the first work to demonstrate that contextual sparsity can scale effectively to large batch sizes, delivering substantial inference acceleration with minimal changes, making Polar Sparsity practical for large-scale, high-throughput LLM deployment systems. Our code is available at: https://github.com/susavlsh10/Polar-Sparsity.
@@ 1 Introduction
(a) Decode latency breakdown
(b) Batch Neuron Activations (Random Sequences)
Figure 1: OPT 66b analysis, A100 GPUs, seq len 1920. (a) LLM Decoding Stage latency; Attention layers dominates at scale. (b) Union neuron sparsity diminishes with batching.
@@
Activation sparsity has been shown to be effective in reducing inference latency, but its application has been largely limited to single-query workloads Dejavu ; powerinfer ; llminaflash . Modern inferencing systems depend on batching to maximize hardware efficiency and reduce serving costs vLLM ; baton ; batchllm ; optimizingllminferencethroughput . Neural scaling laws have driven the rise of massive models that rely on multi-GPU inference, rendering single-query execution prohibitively expensive for real-world deployment scalinglaws ; llama3 ; deepseekv3 . Recent work on activation sparsity focus on reducing latency by optimizing neuron activations in MLP and linear projection layers as these modules dominate runtime in small batch sizes. However, as shown in Figure 1, this strategy breaks down under batched workloads. While batching improves the compute efficiency of linear layers, it has the opposite effect on attention, which scales with sequence length and batch size. Furthermore, union neuron sparsity diminishes with batching (Figure 1), limiting the practical gains of existing approaches. This work explores techniques for integrating contextual sparsity into batched inference, enhancing both system throughput and computational efficiency.
Polar Sparsity refers to the shift in sparsity importance from MLP layers to Attention layers as batch size and sequence length increase, driven by the linear growth of attention layer complexity during the decode stage. Current state-of-the-art sparsity methods primarily focus on model parameter sparsity, where only a subset of model parameters is activated to reduce computation and memory IO. We investigate activation patterns in MLP layers as we scale batch size in naturally sparse models. Figure 1 shows that initial layers display significantly higher sparsity, even with larger batch sizes, indicating that these layers predominantly activate a smaller subset of neurons, providing an opportunity to exploit their inherent sparsity for acceleration.
@@
In large-batch inference, the cost of accessing model parameters is largely amortized, since the entire batch utilizes the same model weights. In contrast, each batch has a unique key-value (KV) cache, making attention layers memory I/O expensive. While contextual sparsity in model parameters diminishes as batch sizes increase, attention head sparsity remains stable and batch invariant. We leverage these properties to build contextual sparsity-aware Selective GEMM and Selective Head/Group FlashAttention kernels that reduce memory I/O and compute, enabling scalable and high-throughput inference. We will release all relevant code, including custom kernels, to encourage reproducibility and future research. To the best of our knowledge, we are the first work to show that contextual sparsity is scalable with batch size, offering even higher gains at larger batch sizes. Our main contributions are as follows:
@@
    We design Selective GEMM kernels with a layer-wise top-$k$ optimization strategy for dynamic MLP activations, achieving up to $\times5$ speedup.
@@
    We introduce Selective FlashAttention kernels that support Head/Group sparsity with per-query activation, reducing memory I/O and compute, achieving up to 2.8×2.8$\times$2.8 × speedup.
@@
    Polar Sparsity delivers up to $\times2$ improvement in batched decoding throughput with negligible accuracy loss, and can be seamlessly integrated into a wide range of LLMs, including those without ReLU activations, to unlock substantial inference acceleration.
@@ 2 Background and Related Works
Several works explore token sparsity in attention, leveraging the fact that many tokens contribute little to model output hashattention ; a2sfaccumulativeattentionscoring ; attentionscoreisnotallyouneed ; keyformer ; NSA ; MoBA . This is orthogonal to head sparsity. MoA MoA and MoH MoH apply MoE-style routing to attention heads and demonstrate gains in accuracy and theoretical efficiency, but these approaches do not lead to wall-clock speedups, indicating that reduced compute does not necessarily result in faster inference. While DejaVu Dejavu and TEAL tealsparsity also exploit head sparsity, their efficiency gains are limited to the projection layers. These methods compute QKV vectors only for selected heads and copy the corresponding KV cache into a smaller buffer before applying dense attention. However, this copying step introduces additional memory I/O, limiting scalability. Moreover, skipping heads during QKV projection can result in missing context for future decoding steps. To overcome these issues, we propose a selective sparsification strategy: we retain dense QKV projections for consistency in KV cache and apply head sparsity only within the attention kernel itself. Additionally, we introduce an I/O-efficient Selective FlashAttention kernel that preserves sparsity benefits while efficiently scaling to larger batch sizes.
Existing methods for improving batch efficiency using contextual sparsity are limited in scope. PowerInfer considers batch sizes up to 32 but targets CPU-based offloading systems, where the baseline is already constrained by CPU and PCIe bandwidth powerinfer . Herd attempts to group sequences with similar sparsity patterns to increase overlap herd , but operates at batch sizes $\leq$4 and is difficult to apply in dynamic, real-world random queries, where identifying on-the-fly sequences that share similar activation patterns becomes prohibitively challenging. These approaches sparsify only the MLP block and speedup decreases with larger batch sizes. In contrast, our method handles arbitrary sequences and scales effectively to large batch sizes, aligning with realistic deployment scenarios.
@@ 3 Motivation and Problem Formulation
The inference process in LLMs consists of two stages: prefill and decode. The prefill stage processes the input prompt in a single forward pass to build and cache the KV representations at each transformer layer, and is typically compute-bound. The decode stage then generates tokens autoregressively using the cached KV data; this stage is generally I/O-bound, as each token requires minimal computation but frequent memory access. In practice, decode dominates end-to-end latency, especially when generating long outputs spanning hundreds or thousands of tokens. This work focuses on optimizing the autoregressive decode stage under batched inference, aiming to improve throughput and efficiency in real-world serving scenarios.
@@ 3.1 Accelerating MLP Layers
where $W_{1,S_{B}},W_{2,S_{B}}\in\mathbb{R}^{d\times|S_{B}|}$ are the dynamically pruned weight matrices corresponding to the active neurons in the batch, and $\sigma$ is the activation function. As batch size increases, the union $S_{B}$ typically grows, reducing the overall sparsity and limiting computational savings. Figure 1 shows the average union activation across transformer layers for different batch sizes in the OPT 66b model during the decode stage. We observe a strong layer-wise trend: early layers maintain high sparsity, while deeper layers exhibit progressively higher union activation. Even at large batch sizes, union activation in early layers remains low as the average per-token activation is under 1%. This trend is encouraging, as it reveals an opportunity to accelerate early layers through sparsity, even in large-scale batching.
@@ 3.2 Accelerating Attention Layers
Figure 2: (a) Perplexity increases gradually when only activating the most important Attention heads/layer. (b) Layer 0 has high importance score across a range of models.
In the decode stage, the input to the attention layer is a batch of inputs $x\in\mathbb{R}^{B\times 1\times d}$, where $B$ is the batch size and $d$ is the model dimension. After applying the QKV projections and reshaping, the query tensor has shape $Q\in\mathbb{R}^{B\times H\times 1\times d_{h}}$, while the key and value tensors are updated and retrieved from the KV cache with shape $K,V\in\mathbb{R}^{B\times H_{kv}\times N\times d_{h}}$, where $H$ is the number of attention heads, $H_{\text{kv}}$ is the number of key-value heads, $d_{h}=d/H$ is the per-head dimension, and $N$ is the current sequence length of each batch  [^1] Some implementations represent the tensors as $Q\in\mathbb{R}^{B\times 1\times H\times d_{h}}$ and $K,V\in\mathbb{R}^{B\times N\times H_{kv}\times d_{h}}$ Each query attends to its own key-value tensor, with self-attention computed independently and in parallel across both heads and batch dimensions.
@@ 4.1 Dynamic Sparsity in MLP Blocks
Our fine-grained MLP router builds on recent advances in sparse inference Dejavu ; powerinfer ; llminaflash , enabling selective activation at the neuron level. Given a hidden state for a batch, $x\in\mathbb{R}^{B\times 1\times d}$, the router is implemented as a lightweight two-layer feedforward network with a bottleneck intermediate layer. Further details are provided in Appendix C.
@@
To efficiently compute sparse MLP activations, we design a custom GPU kernel that fuses indexing and matrix multiplication (Algorithm 3), avoiding the overhead of separate gather-scatter operations. Unlike prior work limited to sparse GEMV, our kernel supports batched GEMM workloads and achieves near-linear speedups with increasing sparsity, with improvements of up to $\times5$ over the dense baseline (Figure 3). Implementation details are provided in Appendix D.
4.2 Stable Sparsity in Attention Blocks
We design our attention routers as single-layer fully connected networks that predict head activations based on attention output norms. The router is trained similarly to the MLP routers: for each input, the top-$k$ attention heads with the highest output norms are considered active and used as supervision targets. The attention router produces scalar logits for each attention head, which are used to select the top-$k$ heads for each batch. Since attention routing is performed independently for each instance, different batches may activate entirely different sets of heads/groups. This results in a batch head index tensor that records the active head indices for each batch.
Given the inputs $\mathbf{Q}\in\mathbb{R}^{B\times H\times 1\times d_{h}}$ and $\mathbf{K,V}\in\mathbb{R}^{B\times H\times N\times d_{h}}$ in HBM, we aim to compute the attention output $\mathbf{O}\in\mathbb{R}^{B\times H\times 1\times d}$ for only the activated attention heads for each batch. Our goal is to reduce the amount of memory access and compute by the factor of induced head sparsity. To achieve this, we fuse the head activation logic into a single GPU kernel and perform selective head attention using a sparsity-aware variant of FlashAttention. During batched decoding, each batch and each head is processed in parallel by a different CUDA thread-block or a Triton program. As outlined in Algorithm 1, we modify the FlashAttention algorithm flashattention to incorporate head sparsity by indexing into the relevant heads during kernel initialization using a batch head index tensor. All memory access logic is updated accordingly to ensure only data from active heads are read from and written to, enabling efficient selective head computation within a unified kernel. Figure 3 shows the forward pass runtime of our Select Head Attention (SHA) kernel compared to dense and standard sparse baselines. The SHA kernel exhibits near-linear speedup with respect to head sparsity, achieving 2.8×2.8$\times$2.8 × improvement at 30% sparsity over the dense baseline. For newer models with group query attention (GQA), we embrace group sparsity as query heads within a group share the KV cache gqa .
@@ 4.2 Stable Sparsity in Attention Blocks
5:$B_{c}=\lfloor M_{SRAM}/(4d)\rfloor$; $(\mathbf{O}_{acc},l_{acc},m_{acc})\leftarrow(\vec{0},0,-\infty)$; $T_{c}=\lceil N_{kv}/B_{c}\rceil$
@@
[…]
@@ 5 Evaluation
Experiment Setting: We evaluate the accuracy of Polar Sparsity compared to dense inference using the OPT, LLaMA-2, and LLaMA-3 model families, considering various model sizes. The evaluation is conducted on nine downstream tasks: COPA copa , OpenBookQA openbookqa , PIQA piqa , RTE RTE , Winogrande winogrande , HellaSwag hellaswag , MMLU MMLU , ARC-easy, ARC-challenging ARC . We utilize the lm-eval-harness framework to measure the accuracy of the models on these benchmarks lmeval . To train the routers, we collected 400,000 token samples from random text sequences extracted from the Wikitext-2 training dataset. All experiments are performed on NVIDIA DGX A100 80GB GPU node servers. For OPT models, we sparsify both MLP and Attention. In contrast, LLaMA models use only attention sparsity, as MLP layers offer limited benefits (Appendix B). We scale each model up to batch size of 512 until it reaches the memory limit. We built on top of the FlashAttention kernel and utilized CUDA graphs to measure the decoding throughput with the included routers.
@@
Table 1: LLM zero-shot evaluation at critical thresholds. Polar Sparsity is competitive with the dense baseline with average accuracy within 1%.
Model COPA OpenBookQA PIQA RTE Winogrande HellaSwag MMLU ARC-E ARC-C Average

### v3

@@ Abstract
Accelerating large language model (LLM) inference is critical for real-world deployments requiring high throughput and low latency. Contextual sparsity, where each token dynamically activates only a small subset of the model parameters, shows promise but does not scale to large batch sizes due to union of active neurons quickly approaching dense computation. We introduce Polar Sparsity, highlighting a key shift in sparsity importance from MLP to Attention layers as we scale batch size and sequence length. While MLP layers become more compute-efficient under batching, their sparsity vanishes. In contrast, attention becomes increasingly more expensive at scale, while their head sparsity remains stable and batch-invariant. We develop Selective Head Attention with hardware-efficient, sparsity-aware GPU kernels, delivering up to $2.2\times$ end-to-end speedups for models like OPT, LLaMA-2 & 3, Qwen, Mistral across various batch sizes and sequence lengths without compromising accuracy. To our knowledge, this is the first work to demonstrate that contextual sparsity can scale effectively to large batch sizes, delivering substantial inference acceleration with minimal changes, making Polar Sparsity practical for large-scale, high-throughput LLM deployment systems. Our code is available at: https://github.com/susavlsh10/Polar-Sparsity.
@@ 1 Introduction
(a) Transformer decode latency breakdown
(b) Selective Head Attention
Figure 1: (a) Decode latency breakdown on A100 GPUs (OPT-66B, sequence length = 1920). As batch size increases, attention layers dominate end-to-end latency. (b) Selective Head Attention only activates the most critical heads for each request and accelerates high throughput inference.
@@
Activation sparsity has been shown to be effective in reducing inference latency, but its application has been largely limited to single-query workloads Dejavu; powerinfer; llminaflash. Modern inferencing systems depend on batching to maximize hardware efficiency and reduce serving costs vLLM; baton; batchllm; optimizingllminferencethroughput. Neural scaling laws have driven the rise of massive models that rely on multi-GPU inference, rendering single-query execution prohibitively expensive for real-world deployment scalinglaws; llama3; deepseekv3. Recent work on activation sparsity focus on reducing latency by optimizing neuron activations in MLP and linear projection layers as these modules dominate runtime in small batch sizes. However, as shown in Figure 1, this strategy breaks down under batched workloads. While batching improves the compute efficiency of linear layers, it has the opposite effect on attention, which scales with sequence length and batch size. Furthermore, union neuron sparsity, which varies dynamically with batches and layers, diminishes with batching (Figure 2), limiting the practical gains of existing approaches. This work explores techniques for integrating contextual sparsity into batched inference, enhancing both system throughput and computational efficiency.
Polar Sparsity refers to the shift in sparsity importance from MLP layers to Attention layers as batch size and sequence length increase. Current state-of-the-art sparsity methods primarily focus on model parameter sparsity, where only a subset of model parameters is activated to reduce computation and memory IO. We define head sparsity as the phenomenon where, for a given token, only a subset of attention heads contribute significantly to the output while the remaining heads have negligible effect. In large-batch inference, the cost of accessing model parameters is largely amortized, since the entire batch utilizes the same model weights. In contrast, each batch has a unique key-value (KV) cache, making attention layers memory I/O expensive. While contextual sparsity in model parameters diminishes as batch sizes increase, attention head sparsity remains stable and batch invariant. We introduce Selective Head Attention (Figure 1), which activates only the most critical heads for each request, preserving overall sparsity and improving compute and memory efficiency.
@@
We leverage these properties to build contextual sparsity-aware Selective GEMM and Selective FlashAttention kernels that reduce memory I/O and compute, enabling scalable and high-throughput inference. To the best of our knowledge, we are the first work to show that contextual sparsity is scalable with batch size, offering even higher gains at larger batch sizes. Our main contributions are as follows:
@@
    We design Selective GEMM kernels with a layer-wise top-$k$ optimization strategy for dynamic MLP activations, achieving up to $5.5\times$ speedup.
@@
    We introduce Selective FlashAttention kernels that support Head/Group sparsity with per-query activation, reducing memory I/O and compute, achieving up to $2.8\times$ speedup.
@@
    Polar Sparsity delivers up to $2.2\times$ improvement in batched decoding throughput with negligible accuracy loss, and can be seamlessly integrated into a wide range of LLMs, including those without ReLU activations, to unlock substantial inference acceleration.
@@ 2 Background and Related Works
Several works explore token sparsity in attention, leveraging the fact that many tokens contribute little to model output hashattention; a2sfaccumulativeattentionscoring; attentionscoreisnotallyouneed; keyformer; NSA; MoBA. This is orthogonal to head sparsity. MoA MoA and MoH MoH introduce MoE-style routing over attention heads, improving accuracy and theoretical efficiency but without wall-clock gains—showing that reduced FLOPs do not always yield faster inference. DejaVu Dejavu and TEAL tealsparsity also exploit head sparsity but only within the projection layers: they compute QKV vectors for selected heads and copy the corresponding KV cache into smaller buffers, incurring costly memory I/O and limited scalability. Moreover, skipping heads during QKV projection can lead to incomplete context in later decoding. To address these issues, we retain dense QKV projections for KV-cache consistency and apply head sparsity within the attention kernel. We further design an I/O-efficient Selective FlashAttention kernel that scales efficiently to large batches while preserving sparsity benefits. Finally, while prior activation sparsity methods target single-query inference, we extend activation sparsity to the large-batch regime via batch-invariant head sparsity and custom GPU kernels, enabling scalable, high-throughput LLM inference.
@@ 3 Motivation and Problem Formulation
The inference process in LLMs consists of two stages: prefill and decode. The prefill stage processes the input prompt in a single forward pass to build and cache the KV representations at each transformer layer, and is typically compute-bound. The decode stage then generates tokens autoregressively using the cached KV data; this stage is generally I/O-bound, as each token requires minimal computation but frequent memory access. While prefill has quadratic complexity in sequence length, it is highly parallelizable across tokens and GPUs; by contrast, the decode stage is sequential, I/O-bound, and dominates wall-clock latency in large batch, long-sequence generation. This work focuses on optimizing the autoregressive decode stage under batched inference, aiming to improve throughput and efficiency in real-world serving scenarios.
@@
(a) OPT 66b batch neuron activations
(b) ReLU Llama-2-7b batch neuron activations
Figure 2: Batch neuron activations in MLP layers in sparse models. Union neuron sparsity in MLP layers decreases rapidly with batch size, illustrating why MLP sparsity is less effective at scale.
@@ 3.1 Accelerating MLP Layers
where $W_{1,S_{B}},W_{2,S_{B}}\in\mathbb{R}^{d\times|S_{B}|}$ are the dynamically pruned weight matrices corresponding to the active neurons in the batch, and $\sigma$ is the activation function. As batch size increases, the union $S_{B}$ typically grows, reducing the overall sparsity and limiting computational savings. Figure 2 shows the average union activation across transformer layers for different batch sizes in the OPT 66b model during the decode stage. We observe a strong layer-wise trend: early layers maintain high sparsity, while deeper layers exhibit progressively higher union activation. Even at large batch sizes, union activation in early layers remains low as the average per-token activation is under 1%. This trend is encouraging, as it reveals an opportunity to accelerate early layers through sparsity, even in large-scale batching. In contrast, sparsity patterns in sparsified LLaMA-2-7B models are more sensitive to batching. Figure 2 shows that replacing SwiGLU with ReLU yields high initial sparsity that quickly diminishes with batching. See Appendix B for additional discussion on MLP sparsity.
@@ 3.2 Accelerating Attention Layers
In the decode stage, the input to the attention layer is a batch of inputs $x\in\mathbb{R}^{B\times 1\times d}$, where $B$ is the batch size and $d$ is the model dimension. After applying the QKV projections and reshaping, the query tensor has shape $Q\in\mathbb{R}^{B\times H\times 1\times d_{h}}$, while the key and value tensors are updated and retrieved from the KV cache with shape $K,V\in\mathbb{R}^{B\times H_{kv}\times N\times d_{h}}$, where $H$ is the number of attention heads, $H_{\text{kv}}$ is the number of key-value heads, $d_{h}=d/H$ is the per-head dimension, and $N$ is the current sequence length of each batch  [^1] Some implementations represent the tensors as $Q\in\mathbb{R}^{B\times 1\times H\times d_{h}}$ and $K,V\in\mathbb{R}^{B\times N\times H_{kv}\times d_{h}}$ . Each query attends to its own key-value tensor, with self-attention computed independently and in parallel across both heads and batch dimensions.
@@
Figure 3: (a) Perplexity increases gradually when only activating the most important attention heads/layer, showing that many heads can be skipped with minimal degradation. (b) Layer 0 has high importance score across a range of models, motivating the use of dense attention in early layers.
@@ 4.1 Dynamic Sparsity in MLP Blocks
Our fine-grained MLP router builds on recent advances in sparse inference Dejavu; powerinfer; llminaflash, enabling selective activation at the neuron level. Given a hidden state for a batch, $x\in\mathbb{R}^{B\times 1\times d}$, the router is implemented as a lightweight two-layer feedforward network with a bottleneck intermediate layer, and we trained it using a binary cross-entropy loss with the AdamW optimizer. Further details are provided in Appendix C.
@@
To efficiently compute sparse MLP activations, we design a custom GPU kernel that fuses indexing and matrix multiplication (Algorithm 3), avoiding the overhead of separate gather-scatter operations. Unlike prior work limited to sparse GEMV, our kernel supports batched GEMM workloads and achieves near-linear speedups with increasing sparsity, with improvements of up to $5.5\times$ over the dense baseline (Figure 4). Implementation details are provided in Appendix D.
4.2 Stable Sparsity with Selective Head Attention
We design our attention routers as single-layer fully connected networks that predict head activations based on attention output norms. The router is trained similarly to the MLP routers: for each input, the top-$k$ attention heads with the highest output norms are considered active and used as supervision targets. The attention router produces scalar logits for each attention head, which are used to select the top-$k$ heads for each batch. Since attention routing is performed independently for each instance, different batches may activate entirely different sets of heads. This results in a batch head index tensor that records the active head indices for each batch. Selective Head Attention (SHA) then uses this tensor to perform attention exclusively on the activated heads within a unified kernel.
@@ 4.2 Stable Sparsity with Selective Head Attention
6:Load $\mathbf{q}\in\mathbb{R}^{1\times d}$ from $\mathbf{Q}[b,\text{head\_idx},0,:]$ $\triangleright$ Get the activated query vector for the batch
@@
13:Write $\mathbf{O}_{acc}$ to $\mathbf{O}[b,\text{head\_idx},0,:]$
Given the inputs $\mathbf{Q}\in\mathbb{R}^{B\times H\times 1\times d_{h}}$ and $\mathbf{K,V}\in\mathbb{R}^{B\times H\times N\times d_{h}}$ in HBM, we aim to compute the attention output $\mathbf{O}\in\mathbb{R}^{B\times H\times 1\times d}$ for only the activated attention heads for each batch. Our goal is to reduce the amount of memory access and compute by the factor of induced head sparsity. To achieve this, we fuse the head activation logic into a single GPU kernel and perform selective head attention using a sparsity-aware variant of FlashAttention. During batched decoding, each batch and each head is processed in parallel by a different CUDA thread-block or a Triton program. As outlined in Algorithm 1, we modify the FlashAttention algorithm flashattention to incorporate head sparsity by indexing into the relevant heads during kernel initialization using a batch head index tensor. All memory access logic is updated accordingly to ensure only data from active heads are read from and written to, enabling efficient selective head computation within a unified kernel. Figure 4 shows the forward pass runtime of our SHA kernel compared to dense and standard sparse baselines. The SHA kernel exhibits near-linear speedup with respect to head sparsity, achieving $2.8\times$ improvement at 30% sparsity over the dense baseline. For newer models with group query attention (GQA), we embrace group sparsity as query heads within a group share the KV cache gqa.
@@ 5 Evaluation
Experiment Setting: We first evaluate Polar Sparsity on pretrained models, including OPT optpaper, LLaMA-2 llama2, LLaMA-3 llama3, Mistral mistral7b across multiple model sizes. The evaluation is conducted on nine downstream tasks: COPA copa, OpenBookQA openbookqa, PIQA piqa, RTE RTE, Winogrande winogrande, HellaSwag hellaswag, MMLU MMLU, ARC-easy, ARC-challenging ARC. We utilize the lm-eval-harness framework to measure the accuracy of the models on these benchmarks lmeval. We further evaluate instruction-tuned models on MMLU-PRO MMLUPRO and LongBench longbench. Specifically, we assess LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct, and Qwen-2.5-14B-Instruct qwen2.5model.
To train the routers, we collected 400,000 token samples from random text sequences extracted from the Wikitext-2 training dataset. All experiments are performed on NVIDIA DGX A100 80GB GPU node servers. Sparsity is applied to both the MLP and attention layers only in OPT models. For all other models, we apply only attention head sparsity, since their MLP layers use non-ReLU activations, making MLP sparsification less effective (Appendix B). We scale each model up to batch size of 512 until it reaches the memory limit. We built on top of the FlashAttention triton kernel and utilized CUDA graphs to measure the decoding throughput with the included routers.
@@
Table 1: LLM zero-shot evaluation at critical thresholds. Polar Sparsity (PS) is competitive with the dense baseline with average accuracy within 1%.
Model COPA OBQA PIQA RTE WG HS MMLU ARC-E ARC-C Average